64a^2-25=0 solve the solution set
Answer by Grinnell(63)

And you thought all was lost!
You are dealing with nice perfect squares here!
(8a-5)(8a+5)=0
8a-5=0, a=5/8 (8a=5)
8a+5=0, a=-5/8 (8a=-5)
Get some rest!!